OpenStudy (anonymous):

Hello, please help me with this, I'm really bad at word problems :( What is the equation using simple equation and one unknown. a Group of 20 people, adults and children, pay $64 to go to a matinee. The entrance fee for adults is $6 for children $2. How many adults in the group?? Thanks.

OpenStudy (anonymous):

\[6(20-y)+2y=64\]

OpenStudy (anonymous):

where y is the children then 20-y are the adults

OpenStudy (aravindg):

intrepret like this : \[\large \bf \text{20 people} \rightarrow \text{adults and children }\] they pay $64 so sum of money paid by adults and children should be $64 for adults charge is $6 and for children $2 Suppose x number of adults were there and y number of children were there so total cost will be \[$6 x+$2y=$64\] we are asked how many adults were there ,and that is x in this equation , so solve for x :)

OpenStudy (anonymous):

in other meaning\[4x+40=64\]

OpenStudy (anonymous):

oh okay. thanks I will study both solutions thanks so much

OpenStudy (anonymous):

oh AracindG I cna only use one unknown :(

OpenStudy (aravindg):

@levinelevine please do not provide direct answer like that .Instead help the asker get to the solution .refer openstudy COC http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok haha sorry... sooooo

OpenStudy (aravindg):

x is number of adults

OpenStudy (aravindg):

that will mean number of adults will depend on only one variable and that is y

OpenStudy (aravindg):

got it @miszery ?

OpenStudy (anonymous):

I dunno, I think for me since word problems are confusing, I find it best to have the answer and study how it got there. I get confuse trying to solve it otherwise. thanks guys

OpenStudy (anonymous):

fisrt you have 2 equations, x+y=20, where x ae the adult and y the children. then you have another equation which is 6x+2y=64. you need to solve for x wich is adults. so first solve for y on the fisrt equation (y=20-x) then input it on the second equation (6x+2(20-x)=64 then simplify and you get 4x+40=64... solve for x and you've got it

OpenStudy (anonymous):

okay I unerstand now, so the other unknown is an equation in itself embedded into the second?

OpenStudy (anonymous):

if x's and y's cause you truble use whole words like instead of x use adult and instead of y use children. trust me its better that way 'cause when you'll do complicated stuff it will be easier

OpenStudy (aravindg):

hey @miszery you can write my x as (20-y)

OpenStudy (anonymous):

yesss there are 2 equations and 2 variables

OpenStudy (aravindg):

so you will get an equation in y

OpenStudy (aravindg):

you can solve for y !

OpenStudy (anonymous):

ok yay! thanks guys will work on this on paper now and study it ... it will take a bit LOL xx cheers